# ExPO: Unlocking Hard Reasoning with Self-Explanation-Guided Reinforcement Learning

### v1

@@ Abstract
Recent advances in large language models have been driven by reinforcement learning (RL)-style post-training, which improves reasoning by optimizing model outputs based on reward or preference signals. These methods typically rely on distinguishing between high- and low-quality outputs, increasing the likelihood of generating positive samples while suppressing that of negative ones. GRPO-style approaches implement this by using self-generated samples labeled by a verifier. However, these methods depend heavily on the model’s initial ability to produce positive samples. They primarily refine what the model already knows (distribution sharpening) rather than enabling the model to solve problems where it initially fails. This limitation is especially problematic in early-stage RL training and on challenging reasoning tasks, where positive samples are unlikely to be generated. To unlock reasoning ability in such settings, the model must explore new reasoning trajectories beyond its current output distribution. Such exploration requires access to sufficiently good positive samples to guide the learning. While expert demonstrations seem like a natural solution, we find that they are often ineffective in RL post-training. Instead, we identify two key properties of effective positive samples: they should (1) be likely under the current policy, and (2) increase the model’s likelihood of predicting the correct answer. Based on these insights, we propose Self-Explanation Policy Optimization (ExPO)—a simple and modular framework that generates such samples by conditioning on the ground-truth answer. ExPO enables efficient exploration and guides the model to produce reasoning trajectories more aligned with its policy than expert-written CoTs, while ensuring higher quality than its own (incorrect) samples. Experiments show that ExPO improves both learning efficiency and final performance on reasoning benchmarks, surpassing expert-demonstration-based methods in challenging settings such as MATH level-5, where the model initially struggles the most. Corresponding to: ruiyang.zhou@utexas.edu, shuozhe.li@utexas.edu, leqiliu@utexas.edu.
@@
Keywords Self-improvement  $\cdot$ Guided exploration  $\cdot$ Reinforcement learning with verifiable rewards  $\cdot$ Reasoning  $\cdot$ Bootstrapping
@@ 3.2 Property 1: In-distribution
[…]
@@ 4.2 Comparing Self-Explanation with Expert CoT and Standard CoT
Figure 2: Negative log-likelihood of self-explanation $\tilde{\bm{c}}$ and expert CoT for Qwen2 3B-Instruct on test split.
@@
Figure 3: Winrate labelled by GPT-4o in terms the number of correct steps of the self-explanation $\tilde{\bm{c}}$ and standard self-generated CoT $\bm{c}$ for Qwen2 3B-Instruct on test split.
@@ 6.2 Results on ExP-GRPO
Figure 4: Accuracy on level-5 questions from the MATH test set for Qwen2.5-3B-Instruct (left) and LLaMA-3.2-3B-Instruct (right) over global training steps. ExP-GRPO consistently outperforms both standard GRPO and GRPO SFT-GT-CoT with supervised fine-tuning on $\bm{c}_{E}$, demonstrating its effectiveness in settings where learning signals are sparse. Shaded regions denote standard deviations over 3 random seeds.
@@
Figure 5: Accuracy on the MATH test set for Qwen2.5-3B-Instruct (left) and LLaMA-3.2-3B-Instruct (right) across global training steps. ExP-GRPO consistently outperforms both GRPO and GRPO SFT-GT-CoT, the latter of which uses supervised fine-tuning on expert CoT $\bm{c}_{E}$. These results demonstrate that ExP-GRPO provides more effective and generalizable learning signals, leading to improved sample efficiency and higher overall performance.
@@ Appendix A Proofs
where the second equality holds because of Bayes’ rule.
@@ C.2 Outputs at Initialization
Full example of direct CoT vs explanation for the question from MATH Question: A box contains six cards. Three of the cards are black on both sides, one card is black on one side and red on the other, and two of the cards are red on both sides. You pick a card uniformly at random from the box and look at a random side. Given that the side you see is red, what is the probability that the other side is red? Express your answer as a common fraction. Direct CoT: To solve this problem, we’ll break it down into steps. Step 1: Count the total number of sides that are red. - There are 2 cards that are red on both sides, so each has 2 red sides. That’s 2 * 2 = 4 red sides. - There is 1 card that is black on one side and red on the other, so it has 1 red side. - Total red sides = 4 + 1 = 5 Step 2: Determine the total number of sides in the box. - There are 3 cards that are black on both sides, so each has 2 sides. That’s 3 * 2 = 6 sides. - There are 2 cards that are red on both sides, so each has 2 sides. That’s 2 * 2 = 4 sides. - Total sides = 6 + 4 = 10 Step 3: Calculate the probability that the other side is red given that the side you see is red. - There are 5 red sides in total, and 5 of those are from cards that are red on both sides. - Since we see a red side, we have 5 options for the side we see, and 5 of those are from cards that are red on both sides. - The probability that the other side is red given that the side you see is red is the number of red sides from cards that are red on both sides divided by the total number of red sides. - Probability = Number of red sides from cards that are red on both sides / Total red sides = 5 / 5 = 1 Therefore, the probability that the other side is red given that the side you see is red is 111. Explanation: To solve this problem, we first need to find the total number of red sides on all the cards in the box. There are two cards that are red on both sides, so they have 4 red sides in total. There is one card that is black on one side and red on the other, so it has 1 red side. In total, there are 4 + 1 = 5 red sides. Now, we need to find the probability that the other side is red given that the side we see is red. This is a conditional probability problem. There are 5 red sides out of a total of 6 sides in the box. So, the probability that we see a red side is 5/6. If we see a red side, it could be from the 4 red sides of the cards that are red on both sides, or from the 1 red side of the card that is black on one side and red on the other. The 4 red sides of the cards that are red on both sides are out of the 5 red sides we see. So, the probability that the other side is red, given that we see a red side, is 4/5. The answer is based on the idea of conditional probability. We know that there are 5 red sides out of a total of 6 sides in the box. When we see a red side, it could be from 4 different cards, each with 2 red sides, or from 1 card with 1 red side. Therefore, the probability that the other side is red, given that we see a red side, is 4 out of the 5 red sides.
C.3 Deceptive Loss Dynamics in DPO Training
C.4 Additional Result for ExP-DPO
@@
Table 4: PASS@4 performance on MATH and GSM8K with models trained with ExP-DPO $\tilde{\bm{c}}_{+}$ and expert CoT $\bm{c}_{E+}$. ExP-DPO consistently outperformed $\bm{c}_{E}$ trained policy in both online and offline setting.
@@
C.5 Additional Result for ExP-GRPO

### v2

@@ Abstract
Self-improvement via RL often fails on complex reasoning tasks because GRPO-style post-training methods rely on the model’s initial ability to generate positive samples. Without guided exploration, these approaches merely reinforce what the model already knows (distribution-sharpening) rather than enabling the model to solve problems where it initially generates no correct solutions. To unlock reasoning ability in such settings, the model must explore new reasoning trajectories beyond its current output distribution. Such exploration requires access to sufficiently good positive samples to guide the learning.
While expert demonstrations seem like a natural solution, we find that they are often ineffective in RL post-training. Instead, we identify two key properties of effective positive samples: they should (1) be likely under the current policy, and (2) increase the model’s likelihood of predicting the correct answer. Based on these insights, we propose Self-Explanation Policy Optimization (ExPO)—a simple and modular framework that generates such samples by conditioning on the ground-truth answer. It can be integrated with popular RL training methods like GRPO and DPO. ExPO enables efficient exploration and guides the model to produce reasoning trajectories more aligned with its policy than expert-written CoTs, while ensuring higher quality than its own (incorrect) samples. Experiments show that ExPO improves both learning efficiency and final performance on reasoning benchmarks, surpassing expert-demonstration-based methods in challenging settings such as MATH level-5, where the model initially struggles the most. Code available in https://github.com/HumainLab/ExPO_rl_reasoning_by_explanation.
@@ 3.2 Property 1: In-distribution
for some weighting $w\in\mathbb{R}$. The weighting depends on the relative quality of the data. When a sample is considered “positive”—that is, it performs better than other samples in the batch (e.g., it achieves a higher objective value or is correct)—the weighting is positive ($w>0$). Conversely, samples that perform worse has the weighting to be w<0w<0.222For certain algorithms (e.g., PPO), the weighting is applied per-token, i.e., the weighting is different for each of the summand in $\nabla_{\theta}\log\pi_{\theta}(\bar{c},\bar{a}|\bar{q})=\sum_{t}\nabla_{\theta}\log\pi_{\theta}(\bar{c}_{t}|\bar{q},\bar{c}_{<t})+\nabla_{\theta}\log\pi_{\theta}(\bar{a}|\bar{q},\bar{c})$. For illustration simplicity, we keep a fixed weighting for each token in a (response) trajectory, but the general properties we illustrate below apply to settings with token-dependent weighting.
@@ 4.2 Comparing Self-Explanation with Expert CoT and Standard CoT
Figure 2: Left: Negative log-likelihood of the self-explanation $\tilde{\bm{c}}$ and expert CoT. Right: Winrate labelled by GPT-4o in terms the number of correct steps of the self-explanation $\tilde{\bm{c}}$ and self-generated CoT $\bm{c}$. Both on the test split of each dataset using Qwen2.5 3B-Instruct.
@@ 6.2 Results on ExP-GRPO
Figure 3: Accuracy on the level-5 questions from MATH test set (left 1, 2) and on the whole test set (right 3, 4) for Qwen2.5-3B-Instruct and LLaMA-3.2-3B-Instruct across global training steps. ExP-GRPO consistently outperforms both GRPO and GRPO SFT-GT-CoT, the latter uses supervised fine-tuning on expert CoT $\bm{c}_{E}$. The results show that ExP-GRPO provides more effective and generalizable learning signals, leading to improved sample efficiency and higher overall performance.
@@ 7 Discussion and Conclusion
Discussion of broader application.
While our experiments focus on math reasoning, ExPO’s core idea of bootstrapping learning with self-generated, in-distribution samples conditioned on verifiable outcomes applies broadly. For example, in code generation (e.g., Codeforces), correct outputs can be verified via test cases. Given the desired output, the model can be prompted to generate rationale (e.g., pseudocode), enabling ExPO to provide effective training signals. In common-sense reasoning tasks where answers are deterministic, ExPO can generate the multi-step reasoning required to reach the correct answer. More generally, any reasoning task with verifiable rewards (such as physics or scientific QA) can benefit from this approach. Ultimately, ExPO demonstrates that a model can effectively teach itself without expert-labeled chain-of-thought, paving the way for enhancing language models’ complex reasoning ability.
Discussoin of the exploratory role of the advantage-weighted objective.
ExPO method uses the relative quality as learning signals of the self-generated explanation and direct chain-of-thought. While Lemma 2 guarantees that the better-but-may-be-imperfect self-explanations often provide partial or heuristic reasoning paths that still guides the model to learn, we want to further discuss why the imperfect CoTs will not make the model collapse. On the one hand, our method does not collapse to blindly imitating these imperfect CoTs. The ExPO training objective includes a reinforcement term with an advantage-weighted update, which continues to explore around these initial CoTs. This ensures that the model does not simply memorize the self-generated explanation $\tilde{c}$, but instead uses it as an anchor for exploration, gradually refining the reasoning policy through trial-and-error. On the other hand, we could employ an annealing schedule for the $\beta$ coefficient: gradually decrease $\beta$ as training progresses. This would make the model rely less on the SFT term in later stages, effectively "fine-tuning" its bootstrapped knowledge during warm-up while reducing the risk of locking in any emergent hallucinations.
@@
Acknolwedgement
This research was supported in part by a research grant from Open Philanthropy. We thank Victor Wang for his valuable discussions and feedback.
@@ Appendix A Proofs
where the second equality holds because of Bayes’ rule. Note that after applying the Bayes rule, $\pi_{\theta}(a^{\star}\mid\tilde{\bm{c}},q)$ cannot be directly generated by the language model, but it is a general distribution (the reverse conditional).
@@ C.2 Outputs at Initialization
The expert-annotated CoT is mathematically dense and concise, demonstrating the solution by expanding and cancelling terms in a single step. In contrast, the self-explanation is more pedagogical and compositional; it first defines the core concept, then explicitly states the key logical insight ($9!=9\times 8!$), and finally performs the simplified calculation. This decomposed, principle-first approach provides a better in-distribution learning signal for RL training.
Example of expert-annotated CoT vs explanation for the question from MATH Question: Compute without using a calculator: $9!/8!$ Expert-annotated CoT: $9!/8!=\frac{9\times 8\times\dots\times 1}{8\times 7\times\dots\times 1}=\boxed{9}$ Self-explanation: The expression $9!/8!$ can be simplified by recognizing that the factorial function $n!$ (n factorial) is defined as the product of all positive integers up to $n$. So $9!=9\times 8\times\dots\times 1$ and $8!=8\times 7\times\dots\times 1$. Therefore, $9!=9\times 8!$. When we divide $9!$ by $8!$, we get: $\frac{9!}{8!}=\frac{9\times 8!}{8!}=9$ Hence, $9!/8!=\boxed{9}$.
We also provide the full example of direct CoT vs explanation for the question from MATH in Subsection 4.1.
@@
Full example of direct CoT vs explanation for the question from MATH Question: A box contains six cards. Three of the cards are black on both sides, one card is black on one side and red on the other, and two of the cards are red on both sides. You pick a card uniformly at random from the box and look at a random side. Given that the side you see is red, what is the probability that the other side is red? Express your answer as a common fraction. Direct CoT: To solve this problem, we’$l$ break it down into steps. Step 1: Count the total number of sides that are red. - There are 2 cards that are red on both sides, so each has 2 red sides. That’s 2 * 2 = 4 red sides. - There is 1 card that is black on one side and red on the other, so it has 1 red side. - Total red sides = 4 + 1 = 5 Step 2: Determine the total number of sides in the box. - There are 3 cards that are black on both sides, so each has 2 sides. That’s 3 * 2 = 6 sides. - There are 2 cards that are red on both sides, so each has 2 sides. That’s 2 * 2 = 4 sides. - Total sides = 6 + 4 = 10 Step 3: Calculate the probability that the other side is red given that the side you see is red. - There are 5 red sides in total, and 5 of those are from cards that are red on both sides. - Since we see a red side, we have 5 options for the side we see, and 5 of those are from cards that are red on both sides. - The probability that the other side is red given that the side you see is red is the number of red sides from cards that are red on both sides divided by the total number of red sides. - Probability = Number of red sides from cards that are red on both sides / Total red sides = 5 / 5 = 1 Therefore, the probability that the other side is red given that the side you see is red is 1. Explanation: To solve this problem, we first need to find the total number of red sides on all the cards in the box. There are two cards that are red on both sides, so they have 4 red sides in total. There is one card that is black on one side and red on the other, so it has 1 red side. In total, there are 4 + 1 = 5 red sides. Now, we need to find the probability that the other side is red given that the side we see is red. This is a conditional probability problem. There are 5 red sides out of a total of 6 sides in the box. So, the probability that we see a red side is 5/6. If we see a red side, it could be from the 4 red sides of the cards that are red on both sides, or from the 1 red side of the card that is black on one side and red on the other. The 4 red sides of the cards that are red on both sides are out of the 5 red sides we see. So, the probability that the other side is red, given that we see a red side, is 4/5. The answer is based on the idea of conditional probability. We know that there are 5 red sides out of a total of 6 sides in the box. When we see a red side, it could be from 4 different cards, each with 2 red sides, or from 1 card with 1 red side. Therefore, the probability that the other side is red, given that we see a red side, is 4 out of the 5 red sides.
C.3 Results for ExP-DPO
@@
We evaluate two variants of ExP-DPO: an offline and an iteratively updated online version. For offline, we first compare the impact of using expert CoTs ($\bm{c}_{E}$) versus self-generated explanations ($\tilde{\bm{c}}$) as the preferred response. As shown in Table 4, training with $\bm{c}_{E}$ leads to only marginal improvements in downstream performance. This observation aligns with our earlier analysis: although expert CoTs are semantically correct, they often lie in low-probability regions under the current policy $\pi_{\theta}$, violating the in-distribution criterion (Property 1). Consequently, they induce misaligned gradient signals. Moreover, from visualization of its loss (see Appendix for details), we observe that even minimal increases in the log-probability of $\bm{c}_{E}$ are sufficient to yield a near-zero loss, creating the illusion of rapid convergence. Therefore, we transition to the online ExP-DPO setting, where the preferred explanation $\tilde{\bm{c}}$ is regenerated periodically from the current policy. These improvements underscore the importance of maintaining distributional alignment between positive training samples and the evolving policy. By continuously adapting $\tilde{\bm{c}}$ to reflect the model’s current beliefs, ExP-DPO ensures that learning remains anchored in the regions where the model can most effectively improve, thereby enabling robust optimization without relying on external expert demonstrations.
@@
C.4 Deceptive Loss Dynamics in DPO Training
C.5 Choice of $\beta$ for ExP-GRPO
In our implementation of ExP-GRPO, the guidance term of self-explanations on the ground-truth answer is incorporated into the overall objective via a scaling coefficient betabeta: $\beta\log\pi_{\theta}(\tilde{\bm{c}},a^{\star}|q)$ This term is designed to provide an additional learning signal when the model’s own sampled responses fail to yield effective gradients, particularly in challenging reasoning settings. We fix $\beta=0.04$ in all reported experiments. To justify this design, we conducted an ablation study across multiple $\beta$ values on the full MATH training set (1 epoch).
Model \ $\beta$ 0.5 0.1 0.08 0.04 0.01 0
Qwen2.5-3B-Instruct 47.3% 48.6% 50.5% 68.7% 60.4% 51.6%
LLaMA-3.2-3B-Instruct 33.5% 34.8% 35.3% 58.9% 46.6% 44.3%
Table 5: Accuracy of ExPO-GRPO with different $\beta$ values on the MATH training set after 1 epoch of training. We can see that $\beta=0.04$ yields to the best training result.
From Table  5, we observe a clear trend: the addition of the ExP-SFT term significantly boosts performance over the baseline ($\beta=0$), particularly in early-stage training where standard GRPO fails to provide meaningful updates. However, excessively large $\beta$ values (e.g., $\beta=0.5$) degrade performance, likely due to over-reliance on imperfect guidance signals. In contrast, smaller $\beta$ values (e.g., $\beta=0.01$) yield slower learning.
Thus, $\beta=0.04$ offers a favorable trade-off between initial learning efficiency and final performance. It balances the model’s ability to leverage the structured supervision from self-generated explanations without overwhelming the exploration driven by the advantage term in GRPO. These results empirically support our design choice and highlight the complementary role of the ExP-SFT term in bootstrapping reasoning capability where standard policy optimization struggles.
C.6 ExPO-GRPO results across different base model sizes
Qwen2.5-1.5B-It LLaMA-3.2-1B-It Qwen2.5-3B-It LLaMA-3.2-3B-It Qwen2.5-7B-It LLaMA-3.1-8B-It
ExP-GRPO 45.2% 40.5% 68.7% 58.9% 83.2% 70.9%
GRPO SFT-GT-CoT 38.8% 35.8% 61.9% 48.2% 79.8% 64.9%
GRPO 32.5% 28.3% 50.3% 44.4% 78.1% 63.9%
Table 6: Accuracy of ExPO-GRPO across different base model sizes on the MATH test set. For >7B model, we use the LoRA parameter-efficient fine-tuning. ExPO consistently outperforms other baselines across all tested scales.
To better demonstrate the generality and robustness of ExPO, we conducted a broader evaluation across model scales. The results show that ExPO performance gains hold even in >=7B models, suggesting that ExPO’s design, which is grounded in in-distribution gradient alignment and positive learning signal, scales effectively with model capacity. These results strengthen our claim that ExPO is a scalable and general reinforcement learning method for reasoning, even in the absence of expert CoT supervision.
C.7 Additional Result for ExP-GRPO